# The Design Space of Lockfiles Across Package Managers

### v1

@@ 6.3. Adoption of Dependency Management Tools within Development Teams
7. Conclusion
We have conducted a comprehensive analysis of lockfiles across 7 package managers, analyzing their documentation and implementation, and collecting quantitative data from open-source GitHub projects and qualitative data from developer interviews. To our knowledge, this is the first in-depth study conducted specifically on lockfiles. We systematically determine how different package managers follow different approaches to generate and enforce lockfiles, and these decisions ultimately result in varied levels of developer satisfaction. We further identify the reasons why developers find lockfiles useful: to achieve deterministic builds, to perform integrity checks, to simplify debugging. We also unveil the unique challenges developers face when using lockfiles: readability issues, indirect dependency locking limitations, and delayed dependency updates. Based on these observations, we provide four recommendations: always generate a lockfile, always commit the lockfile, limit lockfile content to essential metadata, and consider developer experience.
Today, only the Go ecosystem systematically generates and adds lockfiles in version control system. The Go lockfile is particularly readable, concise and clearly separates the list of dependencies from their checksums, making them easier to understand for developers. On the other hand, lockfiles are almost absent from the Java ecosystem. Out of the 7 package managers we study, only Gradle does not generate a lockfile by default. Meanwhile, Maven, the other major package manager for Java does not have a lockfile at all. We command the Maven community add this feature and learn from the best practices to design an informative and usable lockfile.
@@ Acknowledgments
We thank all the interview participants for dedicating their time for this work. We acknowledge the valuable insights provided by Ruy Adorno, a software engineer at vlt. We also thank Ian Arawjo for feedback on early versions of this paper. This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the IVADO consortium.
Appendix A Lockfile Related Functions
npm CLI: npm CLI generates a lockfile by default with the npm install action when executed without additional flags. When no lockfile is available on the first run, npm CLI uses the current state of installed dependencies on disk within the node_modules folder as a starting point. Then, according to the dependency specification files, i.e., package.json, it resolves direct and indirect dependencies by connecting to the npm registry. On the other hand, if a lockfile exists already, npm CLI combines the lockfile and the dependency specification file content to construct the dependency tree. Therefore, the dependency versions resolved would be the same as the locked versions in the lockfile if there are no conflicts between the dependency specification file and the lockfile. If a conflict is detected, npm CLI resolves the dependencies according to the dependency specification file and automatically updates the lockfile. For example, in 1, if the version of the dependency apidevtools/json-schema-ref-parser is updated to ^11.2.0 from ^11.1.0, it conflicts with the version 11.1.1 locked in the lockfile. In such cases, npm CLI resolves a newer available version that satisfies the given version range and updates the lockfile accordingly. When running the npm ci action, the lockfile is validated strictly. This means that npm CLI compares the desired state of dependencies in the dependency specification file, with the dependency tree derived from the lockfile, and raises an error if any mismatch is detected.
pnpm: The pnpm install action functions similar to the npm install action. It creates a lockfile by default when a lockfile does not exist. In case the lockfile exists, it considers the versions locked in the lockfile as long as there are no mismatches with the dependency specification file, or updates the lockfile in case of mismatches. The equivalent behavior to the npm ci action can be achieved with the --frozen-lockfile flag in pnpm.
Cargo: Cargo generates a lockfile by default with cargo build action. However, unlike npm CLI and pnpm, the Cargo build action does not consider the lockfile when resolving dependencies even when one exists. If no lockfile exists or changes to dependencies are required, Cargo automatically writes to the lockfile. To enforce strict use of the locked versions and to disallow writing to the lockfile, developers must enable the --locked or --frozen flags. Compared to the --frozen flag in pnpm, Cargo --frozen flag not only enforces locked versions but also operates in offline mode, supporting air-gapped builds.
Poetry: If no lockfile exists when executing the install or sync actions, Poetry resolves the dependencies and generates a new lockfile by default. If a lockfile exists, Poetry relies on it during the install action, and does not modify it under default configurations. Instead, it raises an error if the lockfile needs to be updated to align with the dependency specification file. During the install action, Poetry compares the content hash of the dependency specification file with the hash recorded in the lockfile’s metadata section. Matching hashes indicates that the dependency specification file has not been updated. However, if the hashes differ due to changes in the dependency specification file, the install process terminates, recommending that the lockfile be updated first by running the poetry update action. The update action resolves the latest versions of dependencies and updates the lockfile accordingly. Similar to npm ci and cargo build –locked actions, Poetry has the poetry sync action that ensures only the dependencies and versions specified in the lockfile are resolved and installed in the project environment.
Pipenv: Pipenv follows a process somewhat similar to Poetry for its install, update, and sync actions. However, under default conditions, Pipenv does not throw an error if the dependency specification file content hash differs from the recorded hash in the lockfile when executing the install action. By default, it also does not strictly rely on the lockfile when resolving dependencies. If new versions are available under the version range defined in the dependency specification file, it will resolve the newer versions irrespective of the locked versions in the lockfile. With Pipenv, the lockfile is only enforced (only the locked dependency versions are considered), and a hash mismatch triggers an error, when the --deploy flag is enabled during the install action. The sync action in Pipenv behaves similarly to Poetry’s, and also enforces locked versions. However, unlike Poetry’s sync action, Pipenv throws an error if a lockfile is missing when the sync action is executed.
Gradle: Gradle takes a distinct approach to generating and enforcing lockfiles compared to other package managers. By default, the Gradle build action neither considers an existing lockfile nor creates a new one when a lockfile does not exist. A lockfile is generated only when the --write-locks flag is enabled. However, even with this flag enabled, a lockfile is not created unless lock states are explicitly defined in the dependency specification file. Developers must specify which groups of dependencies they wish to lock in the dependency specification file. Once these lock states are defined and the --write-locks flag is enabled, Gradle will either create a new lockfile or update an existing one. If a lockfile already exists, Gradle provides options for enforcing the lockfile. In the default and strict modes, Gradle validates the lockfile and throws errors if there are any dependency mismatches. In the lenient mode, however, mismatches with the lockfile do not trigger an error.
Go: Go uses a deterministic dependency resolution algorithm called Minimal Version Selection (MVS), and two files: one called go.mod that works as both a dependency specification file and a lockfile, designed to be machine-written and human-readable, and another called go.sum that stores package checksums. When creating a Go module, the go mod init command generates a go.mod file that contains the Go version and the module path. As they implement their modules, developers import dependencies within the source files that require them. Subsequently, running go mod tidy or any module-building command, such as go build, adds module dependencies resolved according to the import statements to go.mod and generates the go.sum. For additional edits developers can use the go get or go mod edit commands.

### v2

@@ 6.3 Adoption of dependency management tools within development teams
7 Conclusion
We have conducted the first in-depth study of lockfiles across 7 package managers. We analyzed their documentation and implementation, and collected quantitative data from open-source GitHub projects and qualitative data from developer interviews. We systematically documented how different package managers follow different approaches to generate and enforce lockfiles, and these decisions ultimately result in varied levels of developer satisfaction. We further identified the reasons why developers find lockfiles useful: to achieve deterministic builds, to perform integrity checks, to simplify debugging. We also unveiled the unique challenges developers face when using lockfiles: readability issues, indirect dependency locking limitations, and delayed dependency updates. Based on these observations, we provided 2 recommendations for package manager users and 3 recommendations for package manager developers.
Today, only the Go ecosystem systematically generates and adds lockfiles in version control system. The Go lockfile is particularly readable, concise and clearly separates the list of dependencies from their checksums, making them easier to understand for developers. On the other hand, lockfiles are almost absent from the Java ecosystem. Out of the 7 package managers we study, only Gradle does not generate a lockfile by default. Meanwhile, Maven, the other major package manager for Java does not have a lockfile at all. We recommend the Maven community to add this feature and learn from the best practices to design an informative and usable lockfile.
@@ Acknowledgments
We thank all the interview participants for dedicating their time for this work. We acknowledge the valuable insights provided by Ruy Adorno, a software engineer at vlt. We also thank Ian Arawjo for feedback on early versions of this paper. This work is supported by the CHAINS project funded by Swedish Foundation for Strategic Research (SSF), by the Wallenberg Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by IVADO and the Canada First Research Excellence Fund.